The two solid circular bars made of the same linear elastic material are connected by two rigid circular gears Find the magnitude of the torque carried in shaft ac Clearly box your answer rounded to the nearest lb-in 30in 24in GearA rac = 1.00in ?-12 k-in rbd 1.25in T. Gear E r- 2.5 in ??Tac=1362lb-in
Answer #1

Arc length of the gear will be equal. 14c = IBD r04c = r208D 4x IACLAC = 2.5x2.00 G(1PAC Substitute 30 in. for Lac, ( * *2* fAlso, TA_A TA = 1.6T; Here, TBD = Tz Consider the shaft AC. Tac = 12-T. Tac =12–(4.883x1.6T.C) Tac = 12 – 7.8128TAc 8.8128Tc